The economist for the Grand Corporation has estimated the company's cost function, us-

ing time series data, to be

TC = 50 + 16Q -20° + 0.20%

where TC = Total cost I

Q = Quantity produced per period

a. Plot this curve for quantities 1 to 10.

b. Calculate the average total cost. average variable cost, and marginal cost for these quan-

tities, and plot them on another graph.

c. Discuss your results in terms of decreasing, constant, and increasing marginal costs.

Does Grand's cost function illustrate all these?



1
Expert's answer
2021-01-13T10:23:34-0500

"TC = 50 + 16Q - 2Q^2 + 0.2Q^3."

a. This curve for quantities 1 to 10 is upward-sloping.

b. "ATC = TC\/Q = 50\/Q + 16 - 2Q + 0.2Q^2."

"AVC = VC\/Q = 16 - 2Q + 0.2Q^2."

"MC = TC'(Q) = 16 - 4Q + 0.6Q^2."

c. We can observe increasing marginal costs.